OpenStudy (anonymous):

a bad contains a total of 25 marbles 5 of the marbles are blue 3 are red 8 are green 9 are white if a marble is drawn from the bag and then replaced 50 times about how many times would a green marble be drawn A.4 B.8 C.16 D.24

OpenStudy (anonymous):

the ratio of green marbles to the total number of marbles is \(\frac{9}{25}\) your answer is therefore \[\frac{9}{25}\times 50\] cancel first, multiply last

OpenStudy (anonymous):

ooh i read it wrong, it is 8 green sorry \[\frac{8}{25}\times 50\] is the right answer

OpenStudy (anonymous):

Marisol plays a game using the spinner shown. During the game she will spin a total of 54 times. Which is the best prediction for the number of times she will get either "forward 1" or "forward 2" on the spinner?

OpenStudy (anonymous):

please help me thank you very much